OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-4, -3) and (-6, 7). Please show all of your work.

hartnn (hartnn):

can u first find the slope using the equation m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

m=7 - -3/-6 - -4?

hartnn (hartnn):

yes but i would write it as \[m=\frac{7-(-3)}{(-6)-(-4)}\] so what u get m as??

OpenStudy (anonymous):

I have no idea.

hartnn (hartnn):

what is 7-(-3)?

OpenStudy (anonymous):

-4

hartnn (hartnn):

- * - = + so 7-(-3)=7+3=10

OpenStudy (anonymous):

You want me to fill in the -*-?

hartnn (hartnn):

i was telling u that product of 2 negatives is + so here 7-(-3) became 7+3 = 10 did u get this?

OpenStudy (anonymous):

Ooooh ok, got it

hartnn (hartnn):

so now tell me what denominator equals?

OpenStudy (anonymous):

denominator equals -2

hartnn (hartnn):

correct :) so what is m??

OpenStudy (anonymous):

m=-2?

hartnn (hartnn):

m=10/(-2)=-5...see the above formula....ok??

OpenStudy (anonymous):

m= -3/-2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

You are the best

hartnn (hartnn):

so the standard equation is y=mx+c...where u have m now just need to find c you have 2 points for that,choose any 1 say,(-6,7) now u know x,y and m....put it in the equation y=mx+c.....to get c can u?

OpenStudy (anonymous):

ok

hartnn (hartnn):

so give me value of c when u r done...

OpenStudy (anonymous):

hartnn I really appreciate everything you have taught me, I have an emergency here at work and I will have to pick this up when I'm done....I'm so sorry.

hartnn (hartnn):

as u wish ...

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Can someone please start where hartnn left off last night, I had to take care of an emergency at work and could not finish.